# Classical Percolation from Quantum Metric in Flat-Band Delocalization

### 19 Apr 2026 v1

@@ Abstract
The real space quantum metric de Sousa et al. (2023); Marrazzo and Resta (2019); Marsal and Black-Schaffer (2024); Roche Carrasco et al. (2025); Avdoshkin et al. (2025); Romeral et al. (2025) restricted to the occupied states is:
@@
detailed derivations are provided in the supplementary material.
@@
Figure 2: Two terminal transport calculations and geometric conductance simulations. (a) Disorder induced spectral broadening $\eta$ as a function of disorder strength $W$ for both the SOC-free model and the Rashba-SOC model; the dashed horizontal line indicates the band gap $\Delta$. (b) SOC-free case: two-terminal conductance $G$ and geometric conductance $\sigma_{\rm geo}$ versus $W$ for different system sizes $L$; dashed curves denote the total conductance obtained from Green’s function calculation, while symbols show $\sigma_{\rm geo}$ obtained from eq.3. (c) Rashba-SOC case: $G$ and $\sigma_{\rm geo}$ versus $W$ for different $L$, an intermediate diffusive-metal regime emerges. (d) Rashba-SOC case: 2000 sample averaged $\langle\log G\rangle$ versus $W$ for different $L$; inset: finite size scaling near the first crossing yields the critical exponent $\nu$.
@@
In Fig. 2(c), we present the model that includes Rashba SOC. The resulting conductance versus disorder strength curves show a similar trend to that in Fig. 2(b), with flat band localization occurring at $W\leq 2.2$ and Anderson localization at $W\geq 6.5$. However, the conductance values are larger, particularly in the intermediate region where two critical points emerge. In the range $2.2\leq W\leq 6.5$, as the system size increases, the conductance grows, exhibiting metallic behavior. The geometric conductivity calculated using eq.3 closely matches the transport conductance, and the observed larger conductance and metallic behavior can be attributed to the strengthening of the quantum metric due to Rashba SOCSala et al. (2025). To investigate the metallic phase and extract the critical exponent, we calculate the sample averaged log(G)Slevin et al. (2001) and plot it as a function of disorder strength W for various system sizes in Fig. 2(d). The curves reveal two crossing points, signaling a two dimensional inverse Anderson transitionGoda et al. (2006). We fit the critical exponent for the transition from flat band localization to the diffusive metal phase to be approximately 2.24, close to previous studies of 2D symplectic metal-insulator transitionsHikami et al. (1980); Asada et al. (2002, 2004).
To elucidate the microscopic origin of the geometric conductivity observed in the critical window in the absence of SOC, we employ insulating leads by setting the lead Fermi energy to $E_{f}$=0.3, matching the central region, and evaluate the disorder dependence of the sample averaged log(G). In Fig. 3(a), we find a pronounced intermediate regime between flat band localization and Anderson localization where averaged log(G) grows slowly with increasing system size, signaling quasi-metallic behaviorAbrahams et al. (2001). Importantly, because the leads are insulating, this response cannot be attributed to carriers injected from the leads; rather, it reflects spontaneously transport sustained by the intrinsic electronic states of the disordered flat band system, naturally suggesting a percolation type delocalization scenario. Fitting the two critical points $W_{c1}$ and $W_{c2}$ yields $\nu_{1}\sim$1.38 and $\nu_{2}\sim$1.32, close to the two dimensional classical percolation exponent $\nu\sim$1.3Shashkin et al. (1994); Dunford et al. (2001).
Figure 3: Geometric percolation model. (a) SOC-free model with insulating leads: 3000 sample averaged $\langle\log G\rangle$ as a function of disorder strength $W$ for different system sizes $L$. Inset: finite-size scaling analysis in the vicinity of the two crossing points, yielding the critical exponents $\nu_{1}$ and $\nu_{2}$. (b) Schematic of the geometric percolation picture. Colored hemispheres represent effective potential valleys (“puddles”) whose local extent encodes the spread. Neighboring valleys are connected when quantum tunneling is sufficiently strong; in the illustration, the bright links highlight a system-spanning conducting path formed by connected valleys.
To further substantiate this picture, we introduce a geometric percolation modelKirkpatrick (1973); Isichenko (1992); Stauffer and Aharony (1994) rooted in real space quantum metric marker in Fig. 3(b). The gauge invariant Wannier spread $\Omega_{I}$, which quantifies the variance of the real space charge distributionResta and Sorella (1999), can be expressed in terms of a real space quantum metric marker gMarzari and Vanderbilt (1997); Cárdenas-Castillo et al. (2024),
@@
Fig. 4 shows $p_{t}(W)$ for different system sizes $L$. In Fig. 4(a) , around $W\simeq 2$, $p_{t}$ rises sharply from 0 toward $1$ for all $L$, indicating the onset of percolation. This threshold coincides with the transport defined boundary $W_{c1}$ separating the flat-band localized regime from the quasi-metallic window. In the inset of Fig. 4(a), we perform a finite size scaling collapse using the dimensionless scaling variable $z=(W-W_{c})\,L^{1/\nu_{1}}$Dubi et al. (2005a), and find that the data for different $L$ collapse onto a single universal curve, providing strong evidence that the transition is governed by a diverging correlation length with exponent $\nu_{1}$.
In Fig. 4(b), around $W\simeq 4.5$, $p_{t}$ decreases from $1$ back to 0 as $W$ increases, indicating that the system leaves the percolating critical regime and re-enters a localized phase. This second threshold matches the transport defined boundary $W_{c2}$ between the quasi-metallic regime and the Anderson localized regime. The inset of Fig. 4(b) shows an analogous scaling collapse using $z=(W_{c}-W)\,L^{1/\nu_{2}}$ again yielding a single scaling curve. Together, these results support a percolation description of the entire evolution, encoded by the real space quantum metric marker.
@@
The authors would like to thank Ziyue Qi for the helpful discussions. This work was supported by the National Key R&D Program of China (Grant No. 2023YFA1607400, 2022YFA1403800), the National Natural Science Foundation of China (Grant No.12274436, 11925408, 11921004), the Science Center of the National Natural Science Foundation of China (Grant No. 12188101).

### 21 Apr 2026 v2

# Percolation from Quantum Metric in Flat-Band Delocalization

@@ Abstract
The real space quantum metric de Sousa et al. (2023); Marrazzo and Resta (2019); Marsal and Black-Schaffer (2024); Roche Carrasco et al. (2025); Avdoshkin et al. (2025); Mitscherling et al. (2025); Romeral et al. (2025) restricted to the occupied states is:
@@
detailed derivations are provided in the supplementary material. To exclude contribution from the Drude weightKohn (1964); Scalapino et al. (1993); Antebi et al. (2024), we computed, in real space under open boundary conditionsBellomia and Resta (2020), the low frequency Drude weight contribution to the conductivity for different disorder strengths, and found that all of them vanish in the dc limit. The detailed results are presented in the Supplementary Material.
@@
Figure 2: Two terminal transport calculations and geometric conductance simulations. (a) Disorder induced spectral broadening $\eta$ as a function of disorder strength $W$ for both the SOC-free model and the Rashba-SOC model; the dashed horizontal line indicates the band gap $\Delta$. (b) SOC-free case: two-terminal conductance $G$ and geometric conductance $\sigma_{\rm geo}$ versus $W$ for different system sizes $L$; dashed curves denote the total conductance obtained from Green’s function calculation, while symbols show $\sigma_{\rm geo}$ obtained from eq.3. (c) Rashba-SOC case: $G$ and $\sigma_{\rm geo}$ versus $W$ for different $L$, an intermediate diffusive-metal regime emerges. (d) Rashba-SOC case: 1600 sample averaged renormalized localization length versus $W$ for different $L$; inset: finite size scaling near the first crossing yields the critical exponent $\nu$.
@@
In Fig. 2(c), we present the model that includes Rashba SOC. The resulting conductance versus disorder strength curves show a similar trend to that in Fig. 2(b), with flat band localization occurring at $W\leq 2.2$ and Anderson localization at $W\geq 6.5$. However, the conductance values are larger, particularly in the intermediate region where two critical points emerge. In the range $2.2\leq W\leq 6.5$, as the system size increases, the conductance grows, exhibiting metallic behavior. The geometric conductivity calculated using eq.3 closely matches the transport conductance, and the observed larger conductance and metallic behavior can be attributed to the strengthening of the quantum metric due to Rashba SOCSala et al. (2025). To investigate the metallic phase and extract the critical exponent, we calculate the sample averaged renormalized localization lengthMacKinnon and Kramer (1981, 1983); Kramer and MacKinnon (1993); Slevin and Ohtsuki (1999) and plot it as a function of disorder strength W for various system sizes in Fig. 2(d). The curves reveal two crossing points, signaling a two dimensional inverse Anderson transitionGoda et al. (2006). We fit the critical exponent for the transition from flat band localization to the diffusive metal phase to be approximately 2.79, close to previous studies of 2D symplectic metal-insulator transitionsHikami et al. (1980); Asada et al. (2002, 2004).
Figure 3: Schematic of the geometric percolation picture. Colored hemispheres represent effective potential valleys (“puddles”) whose local extent encodes the spread. Neighboring valleys are connected when quantum tunneling is sufficiently strong; in the illustration, the bright links highlight a system-spanning conducting path formed by connected valleys.
To further substantiate this picture, we introduce a geometric percolation modelKirkpatrick (1973); Isichenko (1992); Stauffer and Aharony (1994) rooted in real space quantum metric marker in Fig. 3. The gauge invariant Wannier spread $\Omega_{I}$, which quantifies the variance of the real space charge distributionResta and Sorella (1999), can be expressed in terms of a real space quantum metric marker gMarzari and Vanderbilt (1997); Cárdenas-Castillo et al. (2024),
@@
Fig. 4 shows $p_{t}(W)$ for different system sizes $L$. In Fig. 4(a) , around $W\simeq 2$, $p_{t}$ rises sharply from 0 toward $1$ for all $L$, indicating the onset of percolation. This threshold coincides with the transport defined boundary $W_{c1}$ separating the flat-band localized regime from the quasi-metallic window. In the inset of Fig. 4(a), we perform a finite size scaling collapse using the dimensionless scaling variable $z=(W-W_{c})\,L^{1/\nu_{1}}$Dubi et al. (2005a), and find that the data for different $L$ collapse onto a single universal curve, providing strong evidence that the transition is governed by a diverging correlation length with exponent $\nu_{1}$=1.33, which is classicalShashkin et al. (1994); Dunford et al. (2001).
In Fig. 4(b), around $W\simeq 4.5$, $p_{t}$ decreases from $1$ back to 0 as $W$ increases, indicating that the system leaves the percolating critical regime and re-enters a localized phase. This second threshold matches the transport defined boundary $W_{c2}$ between the quasi-metallic regime and the Anderson localized regime. The inset of Fig. 4(b) shows an analogous scaling collapse using $z=(W_{c}-W)\,L^{1/\nu_{2}}$ with exponent $\nu_{2}$=1.33 again yielding a single scaling curve. Together, these results support a percolation description of the entire evolution, encoded by the real space quantum metric marker.
@@
The authors would like to thank Tobias Holder, Zhida Song, Ziyue Qi for the helpful discussions. This work was supported by the National Key R&D Program of China (Grant No. 2023YFA1607400, 2022YFA1403800), the National Natural Science Foundation of China (Grant No.12274436, 11925408, 11921004), the Science Center of the National Natural Science Foundation of China (Grant No. 12188101).